# ScreenTK: Seamless Detection of Time-Killing Moments Using Continuous Mobile Screen Text and on-device LLM

### v2

@@ Abstract.
Smartphones have become essential to people’s digital lives, providing a continuous stream of information and connectivity. However, this constant flow can lead to moments where users are simply passing time rather than engaging meaningfully. This underscores the importance of developing methods to identify these “time-killing” moments, enabling the delivery of important notifications in a way that minimizes interruptions and enhances user engagement. Recent work has utilized screenshots taken every 5 seconds to detect time-killing activities on smartphones. However, this method often misses to capture phone usage between intervals. We demonstrate that up to 50% of time-killing instances go undetected using screenshots, leading to substantial gaps in understanding user behavior. To address this limitation, we propose a method called ScreenTK that detects time-killing moments by leveraging continuous screen text monitoring and on-device large language models (LLMs). Screen text contains more comprehensive information than screenshots and allows LLMs to summarize detailed phone usage. To verify our framework, we conducted experiments with six participants, capturing 1,034 records of different time-killing moments. Initial results show that our framework outperforms state-of-the-art solutions by 38% in our case study.
@@
ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Figure 1. Comparison between ScreenTK and SOTA screenshot-based method (Chen et al., 2023) for detecting time-killing moments. The green boxes highlight the records of time-killing instances captured by the proposed framework, ScreenTK. Top: passive time-killing study. Bottom: active time-killing study. Transparent: ScreenTK only. Non-Transparent: Both ScreenTK and Screenshot.
Smartphones are essential to modern digital life, providing users with a constant influx of information. However, this constant stream can sometimes be overwhelming, causing users to simply pass time instead of engaging in meaningful activities. This highlights the necessity of developing effective tools to deliver instant and important notifications that help users minimize distractions while enhancing engagement. One effective approach is to identify moments when users are most receptive to incoming content, known as ”attention surplus” moments (Pielot et al., 2015). Within this framework, “time-killing” moments have been identified as a specific type of attention surplus (Chen et al., 2023). Specifically, during “time-killing” moments, users, having no specific goal, seek to fill their perceived free time (Lukoff et al., 2018; Oulasvirta et al., 2012), such as when they are waiting for a train or listening to an unengaging speech (Isaacs et al., 2009).
@@ 3.3. Case Studies
Our goal is to utilize screen text to improve people’s self-awareness and empower self-control over their digital life on smartphones, ultimately benefiting users’ well-being. To achieve this, two case studies were conducted to compare the performance of screen text and screenshots in detecting time-killing behaviors on smartphones. We recruited a total of six volunteer participants from our labs to join the experiment. Data collection was conducted in accordance with ethics approval from our university.
3.4. Passive Time-killing
In the passive test, we aimed to proactively trigger time-killing moments for participants when they were engaged in a reading task. Specifically, a story of 516 words in English was assigned to three participants. After each paragraph, there was a URL embedded in a sentence that read ”Click here.” Additionally, there was a question about the content of the story. The question aimed to prevent participants from merely scanning the content, ensuring they paid more attention to reading the story. To enable a a fair comparison, we used 5-second intervals to capture screenshots and used AWARE-Light to capture screen text information for each user. Data was stored on the device automatically and fed into on-device LLMs when the study was finished.
3.5. Active Time-Killing Study
In the active test, we aim to promote spontaneous time-killing behaviors occurring when participants are focusing on a reading task. Specifically, we designed a reading study using a scientific essay of 2,351 words. The URLs of popular internet memes were embedded in 20 citation brackets. Three participants joined this study without being notified of anything they should be aware of. Other configurations followed those used in the active time-killing text consistently.
@@ 4. Results and Discussion
The passive time-killing study collected 535 records and 11 minutes of screen text from another 3 participants (on average, 178 records and 4 minutes per participant). In comparison, the active time-killing study collected 499 records of screen text from three participants (on average, 11 minutes and 100 records per participant). The quantitative results are discussed in Section 4.1, and the qualitative results are provided in Section 4.2.
@@ 4.1. Quantitative Results
We compared the performance of the proposed ScreenTK by comparing it to the SOTA screenshots method, calculating the capture rate of time-killing actions during the implicit and explicit tests. We observed that the baseline method of taking screenshots every 5 seconds missed many time-killing instances (38% for active time-killing and 57% for passive time-killing). On the other hand, ScreenTK captured almost all time-killing events (except for one instance due to a screen text sensor unresponsive issue). This result suggests that the proposed ScreenTK is significantly more effective in capturing time-killing behaviors compared to the traditional screenshot method.

### v3

@@ Abstract.
Smartphones have become essential to people’s digital lives, providing a continuous stream of information and connectivity. However, this constant flow often depletes users’ limited attentional resources and time, leading to decreased productivity and increased stress levels. This issue underscores the need for tools that empowers users to maximize their potential for achieving personal objectives. One effective approach is to identify “time-killing” moments—a specific type of attention surplus—during which users seek to fill perceived free time without a specific purpose. Recent work has utilized screenshots taken every 5 seconds to detect time-killing activities on smartphones. However, this method often misses to capture phone usage between intervals. We demonstrate that up to 50% of time-killing instances go undetected using screenshots, leading to substantial gaps in understanding user behavior. To address this limitation, we propose a method called ScreenTK that detects time-killing moments by leveraging continuous screen text monitoring and on-device large language models (LLMs). Screen text contains more comprehensive information than screenshots and allows LLMs to summarize detailed phone usage. To verify our framework, we conducted experiments with six participants, capturing 1,034 records of different time-killing moments. Initial results show that our framework outperforms state-of-the-art solutions by 38% in our case study.
@@
journalyear: 2024copyright: rightsretainedconference: Companion of the 2024 ACM International Joint Conference on Pervasive and Ubiquitous Computing Pervasive and Ubiquitous Computing; October 5–9, 2024; Melbourne, VIC, Australiabooktitle: Companion of the 2024 ACM International Joint Conference on Pervasive and Ubiquitous Computing Pervasive and Ubiquitous Computing (UbiComp Companion ’24), October 5–9, 2024, Melbourne, VIC, Australiadoi: 10.1145/3675094.3677547isbn: 979-8-4007-1058-2/24/10submissionid: ubide1153ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Figure 1. Comparison between ScreenTK and SOTA screenshot-based method (Chen et al., 2023) for detecting time-killing moments. The green boxes highlight the records of time-killing instances captured by the proposed framework, ScreenTK. Top: explicit time-killing study. Bottom: implicit time-killing study. Transparent: ScreenTK only. Non-Transparent: Both ScreenTK and Screenshot.
Smartphones are essential to modern digital life, providing users with a constant influx of information. However, this nonstop stream can sometimes be overwhelming, depleting users’ limited attentional resources and hindering the pursuit of personal goals (e.g., learning a language and reading a book). This challenge highlights the need for effective tools that offer users the choice to detect opportune moments to maximize their potential for achieving these objectives. One promising approach is to identify moments when users are most receptive to incoming content, known as ”attention surplus” moments (Pielot et al., 2015). Within this framework, “time-killing” moments have been identified as a specific type of attention surplus (Chen et al., 2023). Specifically, during “time-killing” moments, users, having no specific goal, seek to fill their perceived free time (Lukoff et al., 2018; Oulasvirta et al., 2012), such as when they are waiting for a train or listening to an unengaging speech (Isaacs et al., 2009).
@@ 3.3. Case Studies
Our goal is to utilize screen text to improve people’s self-awareness and empower self-control over their digital life on smartphones, ultimately benefiting users’ well-being. To achieve this, two case studies were conducted to compare the performance of screen text and screenshots in detecting time-killing behaviors on smartphones. We recruited a total of six volunteer participants from our labs to join the experiment (different 3 participants for each case study). The explicit study aims to detect the time-killing moments that are intentionally triggered (Section 3.4) and the implicit study focuses on the time-killing moments that are spontaneous (Section 3.5). Data collection was conducted in accordance with ethics approval from our university.
3.4. Explicit Time-killing Study
In the explicit study, we aimed to proactively trigger time-killing moments for participants when they were engaged in a reading task. Specifically, a story of 516 words in English was assigned to three participants. After each paragraph, there was a URL embedded in a sentence that read ”Click here.” Additionally, there was a question about the content of the story. The question aimed to prevent participants from merely scanning the content, ensuring they paid more attention to reading the story. To enable a a fair comparison, we used 5-second intervals to capture screenshots and used AWARE-Light to capture screen text information for each user. Data was stored on the device automatically and fed into on-device LLMs when the study was finished.
3.5. Implicit Time-Killing Study
In the implicit study, we aim to promote spontaneous time-killing behaviors occurring when participants are focusing on a reading task. Specifically, we designed a reading study using a scientific essay of 2,351 words. The URLs of popular internet memes were embedded in 20 citation brackets. Three participants joined this study without being notified of anything they should be aware of. Other configurations followed those used in the explicit time-killing text consistently.
@@ 4. Results and Discussion
The explicit time-killing study collected 535 records and 11 minutes of screen text from another 3 participants (on average, 178 records and 4 minutes per participant). In comparison, the implicit time-killing study collected 499 records of screen text from three participants (on average, 11 minutes and 100 records per participant). The quantitative results are discussed in Section 4.1, and the qualitative results are provided in Section 4.2.
@@ 4.1. Quantitative Results
We compared the performance of the proposed ScreenTK by comparing it to the SOTA screenshots method, calculating the capture rate of time-killing actions during the explicit and implicit tests. The ground truth of time-killing instances were manually counted by the first author from the sequences of one-second-interval screenshots that were extracted from the screen recordings of the participants’ phone usage. For the explicit and implicit tests, a time-killing instance started when a participant clicked an URL in the story or article and ended when the participant returned back to the original reading. We observed that the baseline method of taking screenshots every 5 seconds missed many time-killing instances (57% for explicit time-killing and 38% for implicit time-killing). On the other hand, ScreenTK captured almost all time-killing events (except for one instance due to a screen text sensor unresponsive issue). This result suggests that the proposed ScreenTK is significantly more effective in capturing time-killing behaviors compared to the traditional screenshot method.
@@ 5. Conclusion and Future Work
Acknowledgements.
This work is partially funded by NHMRC grants 1170937 and 2004316, AUSMURI grant 13203896, and the CISCO grant 2021-2327463696.